OpenStudy (anonymous):

4 fourth root of 5 end root plus 7 cubed root of 5 end root minus 9 cubed root of 5 end root plus 6 fourth root of 5

OpenStudy (anonymous):

can you rephrase the question, maybe using the equation tools :)

OpenStudy (anonymous):

\[4^4\sqrt{5} + 7^3\sqrt{5} - 9^3\sqrt{5} + 6^4\sqrt{5}\] ???

OpenStudy (anonymous):

\[4^{4}\sqrt{5}+ 7^{3}\sqrt{5} - 9^{3}\sqrt{5} + 6^{4}\sqrt{5}\]

OpenStudy (anonymous):

factor out the root 5

OpenStudy (anonymous):

wouldn't all the 5's turn into 1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

think of root 5 as any other variable... X or Y or something and factor it out of there, to get something like: root 5 ( + - + )

OpenStudy (anonymous):

think of 4x + 7x - 9x + 6x

OpenStudy (anonymous):

so factor those out?

OpenStudy (anonymous):

factor out the root 5, that will make it easier... but all you need to do is type it all into a calculator it is just a question of how many buttons you are willing to push

OpenStudy (anonymous):

\[\sqrt{5}(4^4+7^3-9^3+6^4)\]

OpenStudy (anonymous):

so the answer is \[1166\sqrt{5}\]?

OpenStudy (anonymous):

is it 8 sqrt 5?

OpenStudy (anonymous):

I don't think it is 8... these are some big numbers

OpenStudy (anonymous):

looks like josh's calculator has good batteries, I like his answer

OpenStudy (anonymous):

cool, i learnt something new :) thanks philips13

OpenStudy (anonymous):

but that's not of my choices...

OpenStudy (anonymous):

did we copy the problem correctly?